### v3

@@ I Introduction
The DJI, which has a share of more than 70% of the global drone market, encrypts and releases very many kinds of drone firmware with its own code signature. So, it is difficult for firmware analysis tools to recognize and analyze the instruction codes in the DJI drone’s firmware. To avoid these problems, we propose a method for automatic dynamic analysis of increasingly complex and large drone firmware, and the challenges and solutions that may exist. In doing so, we aim to expand research into dynamic analysis of drone firmware and improve security in this area.
@@ I-A OUR APPROACH
Particularly, we employ AFL (American Fuzzy Lop) for fuzzing techniques to effectively identify potential security vulnerabilities within the binary code of drone firmware. This step is crucial for enhancing the security of drone firmware and is expected to significantly contribute to the safe operation of drones.
Additionally, to facilitate knowledge sharing within the research community, we have made the tools we developed open-source. This allows other researchers to easily apply our methodology to their own studies, lowering the barriers to entry for drone firmware security research and aiming to cultivate a safer drone technology ecosystem. This research provides a robust foundation for innovatively advancing dynamic analysis and security research on drone firmware.
I-B CONTRIBUTION
In summary, our takeaways are as follows:
@@ I-B CONTRIBUTION
• How to set up your operating system for dynamic analysis of drone firmware, including the configuration of dependent libraries and the selection of the appropriate operating system version.
@@ III-A Challenges of drone firmware analysis
Operating environment issues with existing analysis tools. For the collected firmware, we used emulators designed for existing firmware analysis or installed existing binary dynamic analysis tools such as GDB server and fuzzer directly on the drone system. However, the emulator or analysis tool designed for existing firmware analysis could not be used due to issues such as dependencies in the environment and firmware libraries and differences in the operating environment. In addition, the method of installing dynamic analysis tools directly on the drone system was successful [13], but there is the problem of not having enough time for dynamic analysis due to limited battery capacity. Drones are designed to operate on battery power, which results in a limited operating time. This limitation makes it difficult to perform comprehensive and time-consuming dynamic analysis on specific programs within the given time frame. Even if the drone system has uninterrupted power supply, the constrained hardware environment, including limited memory and processing power, hinders the efficient execution of dynamic analysis tools, which affects the effectiveness of the analysis. We also used AFL to dynamically analyze the drone firmware. However, we encountered a message from AFL that said, ”Due to a system problem, you can’t run this program (AFL) in this environment. When we analyzed this error, we found that AFL’s code contains features that are not available on the drone’s operating system (Android).
@@ III-B1 Acquiring the binary file for analysis
To decrypt the DJI firmware, you need information about the encryption method used in the firmware and the decryption key. To obtain this information, we first need to analyze the structure of the firmware file to identify the areas where hints about the encryption scheme and decryption key are located. The binary header of the firmware is usually located at the beginning of the file and contains information such as the type, version, and date of the firmware. Most IoT products retrieve information from this binary header to recommend a firmware upgrade to the user or to determine if the firmware is compatible. However, DJI’s firmware has its own signature and encryption for enhanced security, and decryption is performed by reading the header information from a decryption module mounted on the drone’s hardware chip[24]. Therefore, we analyzed the relevant research to collect the header information that DJI drones need to run the firmware[25]. According to the literature on DJI firmware, the DJI drone firmware header contains information about the magic number, version, authentication key identifier, scramble key, etc. Inside the drone, the header information of the corresponding firmware binary is used to verify and decrypt the signature value. The entire DJI drone firmware header is organized as follows, Table 3.
@@ III-B3 Installing dependent libraries
This section describes building an emulation environment for dynamic analysis of decrypted firmware. Firmware analysis involves the use of tools such as emulators and debuggers, and the environment must be configured so that the commands in the analysis tool work properly. Since firmware is highly hardware dependent, the same operating environment as that used by the system to be analyzed must be configured to run the analysis tool. The DJI drone system uses the Android operating system, and the firmware binary does not have PIE (Position Independent Executables), which is mandatory for security reasons in Android 5.0 (Lollipop) and later. Therefore, to build an emulation environment for DJI drone firmware binaries without PIE, an Android 5. In addition, since the DJI drone system uses a customized Android environment, it requires libraries that are used independently. Therefore, we need to check the files used in addition to the libraries used in the Android environment and identify the dependency relationships for each file to the required libraries. We used the command ”ldd” command to identify the required libraries and dependencies, and checked the required files for each binary, as shown in Figure 9. We then, we accessed the internal system of the DJI drone using ADB, extracted the required library files, and installed them on the embedded board to complete the environment setup for dynamic analysis.
@@ III-B4 Installing fuzzer
To undertake an automated analysis of DJI firmware, it is imperative to employ tools like AEG and fuzzer. American Fuzzy Lop (AFL), widely adopted for such purposes, compiles source code via afl-gcc for fuzzing, or resorts to black box fuzzing through user space emulation for binary code. However, given the DJI firmware’s source code is not open to the public and thus unobtainable, analysis is confined to the firmware’s binary form. Upon examining the decrypted DJI firmware, it became apparent that the majority of the binaries are ARM architecture-based, necessitating the use of tools compatible with various architectures, including ARM, Aarch64, and PowerPC. Considering the focus on firmware, vulnerability analysis tools that necessitate source code, such as honggfuzz[27], KLEE[28], and LibFuzzer[29], are not applicable. Consequently, suitable tools for this research include AFL, Radamsa[30], and angr[31], all of which support binaries across different platforms and architectures. Radamsa, however, falls short for this project since it modifies the structure of existing binaries and relies on manual post-execution error checks and analyses. Angr also presents challenges as it requires setting up a simulation manager for each binary to determine probable vulnerability targets, rendering it less viable for extensive automated analysis. Thus, AFL was chosen, a cross-platform fuzzer, alongside an AFL-equipped embedded board, to conduct the firmware’s automated analysis. AFL’s shared memory capabilities facilitate code coverage tracking, performance optimization, and novel input identification. However, the shared memory API it typically uses is restricted to the Android operating systems running drone systems, requiring an alternative API. Because AFL is primarily designed for Linux or Unix-based systems, it is incompatible with Android-based embedded boards. AFL’s reliance on system-specific libraries and debugging tools means that the embedded AFL binary cannot run on different operating systems without modifications that take into account the unique features and constraints of the Android platform. To overcome these shared memory API challenges, an Android-specific AFL version, Android-AFL[32], was discovered and subsequently deployed within the Beaglebone Black embedded board environment. Tailored to the board’s ARM Cortex-A8 processor, a cross-compilation environment was built and configured for this precise architecture. Thus, we adapted Android-AFL for use on embedded boards and built the platform to perform dynamic analysis in a separate analysis environment, as shown in Figure 11.
@@ IV-C Differences between existing research and our proposed method (RQ2)
V Conclusion and Future
We are by addressing the challenges associated with dynamic analysis of drone firmware, particularly that of DJI drones, we have introduced a comprehensive methodology that not only streamlines firmware decryption, but also facilitates automated vulnerability analysis. The creation of a dedicated embedded board, along with the development of an automated decryption module, represents a significant advancement in the field. This setup enables the efficient use of automated analysis tools, overcoming the limitations of traditional methods, which often rely on manual processes and are time-consuming. Our approach improves the efficiency of vulnerability analysis in drone firmware by automating the decryption process for encrypted firmware files and building a drone simulation environment for dynamic analysis will contribute to the advancement of drone security research. Given the increasing reliance on drones in various fields and their potential security vulnerabilities, we provide an automated vulnerability analysis methodology that allows for a more thorough and rapid security assessment of drone firmware. In addition, the methodology and tools developed through this research provide a replicable model for future research that can be extended not only to DJI drones, but also to other UAV systems using the Android environment. This versatility increases the broader applicability of our findings and contributes to the development of safer drone technologies in general. Based on these contributions, it is clear that our work is a critical step towards solving the complex problem of drone firmware security. The automated process introduced here significantly reduces the barriers to comprehensive security analysis, enabling more effective identification and mitigation of vulnerabilities in drone systems.

### v4

@@ I Introduction
The DJI, which has a share of more than 70% of the global drone market, encrypts and releases very many kinds of drone firmware with its own code signature. So, it is difficult for firmware analysis tools to recognize and analyze the instruction codes in the DJI drone’s firmware. To avoid these problems, we propose a method for automatic dynamic analysis of increasingly complex and large drone firmware, and the challenges and solutions that may exist. In doing so, we aim to expand research into dynamic analysis of drone firmware and improve security in this area. In particular, we identified three major challenges:
  • Firmware acquisition and decryption: How to acquiring the drone firmware and making it analyzable?
  • Operating environment issues with existing analysis tools: Drone system has poor performance to perform on-board dynamic analysis
  • Library dependency issues: What are the necessary components for the construction of the dynamic analysis environment for drone firmware?
A detailed description of each challenge is provided in 3.A.
@@ I-A OUR APPROACH
Particularly, we employ AFL (American Fuzzy Lop) for fuzzing techniques to effectively identify potential security vulnerabilities within the binary code of drone firmware. AFL is automated tool to generating and testing of input cases significantly increases test productivity. This automation, combined with its open source nature, allows for continuous improvement and adaptability to different test environments. In addition, the tool’s ability to reproduce crashes and log critical test cases facilitates efficient debugging and problem resolution. AFL is s proven track record, combined with its advanced features and community support, make it an ideal choice for fuzz testing in our research[45]. This step is crucial for enhancing the security of drone firmware and is expected to significantly contribute to the safe operation of drones.
Additionally, to facilitate knowledge sharing within the research community, we have made the tools we developed open-source. This allows other researchers to easily apply our methodology to their own studies, lowering the barriers to entry for drone firmware security research and aiming to cultivate a safer drone technology ecosystem. This research provides a robust foundation for innovatively advancing dynamic analysis and security research on drone firmware. In summary, our contributions to meeting the challenges outlined above are as follows:
@@
  • Set up an operating system for dynamic analysis of drone firmware, including manual steps such as finding dependent libraries and selecting the appropriate operating system version.
@@ II-A Drone Vulnerability Analysis
Proposed works Firmware Dynamic Analysis (Firware binary) Using emulation and fuzzing techniques, we are researching how to perform binary dynamic analysis on drone firmware.
@@ III-A Challenges of drone firmware analysis
Operating environment issues with existing analysis tools. For the collected firmware, we used emulators designed for existing firmware analysis or installed existing binary dynamic analysis tools such as GDB server and fuzzer directly on the drone system. However, the emulator or analysis tool designed for existing firmware analysis could not be used due to issues such as dependencies in the environment and firmware libraries and differences in the operating environment. In addition, the method of installing dynamic analysis tools directly on the drone system was successful [13], but there is the problem of not having enough time for dynamic analysis due to limited battery capacity. Drones are designed to operate on battery power, which results in a limited operating time. This limitation makes it difficult to perform comprehensive and time-consuming dynamic analysis on specific programs within the given time frame. Even if the drone system has uninterrupted power supply, the constrained hardware environment, including limited memory and processing power, hinders the efficient execution of dynamic analysis tools, which affects the effectiveness of the analysis. At this time, it is also possible to consider using multiple sampling, which can obtain more efficient results by sampling various input data several times. However, tools such as MULTIFUZZ[46], PosFuzz[47], Honggfuzz[48], and libFuzzer[49], which are fuzzers using the multiple sampling technique, were not available due to problems such as operating environment, as a result of using them in this study. Furthermore, multiple sampling cannot be effectively applied in situations where CPU and memory limitations exist. Furthermore, the battery usage speed could be accelerated, which made it challenging to utilise the multiple sampling technique. This technique is more suitable for dynamic analysis in an emulation environment, and it is more suitable to apply to tools in emulators or software than drone hardware.
@@ III-B1 Acquiring the binary file for analysis
To decrypt the DJI firmware, you need information about the encryption method used in the firmware and the decryption key. To obtain this information, we first need to analyze the structure of the firmware file to identify the areas where hints about the encryption scheme and decryption key are located.
The binary header of the firmware is usually located at the beginning of the file and contains information such as the type, version, and date of the firmware. Most IoT products retrieve information from this binary header to recommend a firmware upgrade to the user or to determine if the firmware is compatible. However, DJI’s firmware has its own signature and encryption for enhanced security.
Decryption is performed by reading the header information from a decryption module mounted on the drone’s hardware chip[24]. Therefore, we analyzed the relevant research to collect the header information that DJI drones need to run the firmware[25].
According to the literature on DJI firmware, the DJI drone firmware header contains information about the magic number, version, authentication key identifier, scramble key, etc. Inside the drone, the header information of the corresponding firmware binary is used to verify and decrypt the signature value. The entire DJI drone firmware header is organized as follows in Table 3.
@@ III-B3 Installing dependent libraries
This section describes building an emulation environment for dynamic analysis of decrypted firmware. Firmware analysis involves the use of tools such as emulators and debuggers, and the environment must be configured so that the commands in the analysis tool work properly. Since firmware is highly hardware dependent, the same operating environment as that used by the system to be analyzed must be configured to run the analysis tool. The DJI drone system uses the Android operating system, and the firmware binary does not have PIE (Position Independent Executables), which is mandatory for security reasons in Android 5.0 (Lollipop) and later. Therefore, we need to use an Android version lower than 5.0 to build an emulation environment for DJI drone firmware binaries.
In addition, since the DJI drone system uses a customized Android environment, it requires libraries that are used independently. Therefore, we need to check the files used in addition to the libraries used in the Android environment and identify the dependency relationships for each file to the required libraries. We used the command ”ldd” command to identify the required libraries and dependencies, and checked the required files for each binary, as shown in Figure 9. We then, we accessed the internal system of the DJI drone using ADB, extracted the required library files, and installed them on the embedded board to complete the environment setup for dynamic analysis.
@@ III-B4 Installing fuzzer
To undertake an automated analysis of DJI firmware, it is imperative to employ tools like AEG and fuzzer. American Fuzzy Lop (AFL), widely adopted for such purposes, compiles source code via afl-gcc for fuzzing, or resorts to black box fuzzing through user space emulation for binary code. However, given the DJI firmware’s source code is not open to the public and thus unobtainable, analysis is confined to the firmware’s binary form.
Upon examining the decrypted DJI firmware, it became apparent that the majority of the binaries are ARM architecture-based. This necessitates the use of tools compatible with various architectures, including ARM, Aarch64, and PowerPC. Considering the focus on firmware, vulnerability analysis tools that necessitate source code, such as honggfuzz[27], KLEE[28], and LibFuzzer[29], are not applicable. Consequently, suitable tools for this research include AFL, Radamsa[30], and angr[31], all of which support binaries across different platforms and architectures. Radamsa, however, falls short for this project since it modifies the structure of existing binaries and relies on manual post-execution error checks and analyses. Angr also presents challenges as it requires setting up a simulation manager for each binary to determine probable vulnerability targets, rendering it less viable for extensive automated analysis.
AFL was chosen, a cross-platform fuzzer, alongside an AFL-equipped embedded board, to conduct the firmware’s automated analysis. AFL’s shared memory capabilities facilitate code coverage tracking, performance optimization, and novel input identification. However, the shared memory API it typically uses is restricted to the Android operating systems running drone systems, requiring an alternative API. Because AFL is primarily designed for Linux or Unix-based systems, it is incompatible with Android-based embedded boards. AFL’s reliance on system-specific libraries and debugging tools means that the embedded AFL binary cannot run on different operating systems without modifications that take into account the unique features and constraints of the Android platform. To overcome these shared memory API challenges, an Android-specific AFL version, Android-AFL[32], was discovered and subsequently deployed within the Beaglebone Black embedded board environment. Tailored to the board’s ARM Cortex-A8 processor, a cross-compilation environment was built and configured for this precise architecture. Thus, we adapted Android-AFL for use on embedded boards and built the platform to perform dynamic analysis in a separate analysis environment, as shown in Figure 11.
@@ IV-C Differences between existing research and our proposed method (RQ2)
V Conclusion and Future Work
We are by addressing the challenges associated with dynamic analysis of drone firmware, particularly that of DJI drones, we have introduced a comprehensive methodology that not only streamlines firmware decryption, but also facilitates automated vulnerability analysis. The creation of a dedicated embedded board, along with the development of an automated decryption module, represents a significant advancement in the field. This setup enables the efficient use of automated analysis tools, overcoming the limitations of traditional methods, which often rely on manual processes and are time-consuming.
Our approach improves the efficiency of vulnerability analysis in drone firmware by automating the decryption process for encrypted firmware files and building a drone simulation environment for dynamic analysis will contribute to the advancement of drone security research. Given the increasing reliance on drones in various fields and their potential security vulnerabilities, we provide an automated vulnerability analysis methodology that allows for a more thorough and rapid security assessment of drone firmware. In addition, the methodology and tools developed through this research provide a replicable model for future research that can be extended not only to DJI drones, but also to other UAV systems using the Android environment. This versatility increases the broader applicability of our findings and contributes to the development of safer drone technologies in general. Based on these contributions, it is clear that our work is a critical step towards solving the complex problem of drone firmware security. The automated process introduced here significantly reduces the barriers to comprehensive security analysis, enabling more effective identification and mitigation of vulnerabilities in drone systems.